OpenStudy (taylormclauren):

which of the following could be the slope field for the differential equation dy/dx=x^2y

OpenStudy (bobo-i-bo):

You can solve the differential equation since is separable: https://en.wikipedia.org/wiki/Examples_of_differential_equations#Separable_first-order_ordinary_differential_equations

OpenStudy (taylormclauren):

I still don't get it... @Bobo-i-bo

OpenStudy (taylormclauren):

@The_Real_Rapper

OpenStudy (the_real_rapper):

graph iv

OpenStudy (bobo-i-bo):

\[\frac{ dy }{ dx } = x^2y\] So by seperating the variables, using bad maths we can say: \[\frac{ 1 }{ y }dy=x^2dx\] So: \[\int\limits \frac 1 y dy = \int\limits x^2 dx\] And so if you solve the integrals on both sides and rearrange, you can get y as a function of x

OpenStudy (kainui):

When you were in algebra you learned \(y=mx+b\) and m was the slope. Rise over run, right? \[m=\frac{\Delta y}{\Delta x}\] similarly here we have \[\frac{dy}{dx} = x^2y\] so you can imagine this representing the slope at every point in the xy plane. at the point \((-1, 2)\) you can plug this in to get: \[\frac{dy}{dx} = (-1)^22 = 2\] so you find the point on the grid (-1,2) and write a tiny mark with the same slope as the line with m=2 would have.  See how right at that point I made a little line where it looks like it has m=2? Remember, rise over run, so \(m=\frac{2}{1}\) means rise by 2 run to the right by 1. You just want this slope only at this point. Repeat this process for a handful of points and you should get one of the pictures!

OpenStudy (taylormclauren):

So would it be graph III? @Kainui

OpenStudy (the_real_rapper):

yes i understood it wrong yes graph III